# KiDS-Legacy: Redshift distributions and their calibration

### v1

@@ Acknowledgements.
AHW and HHi are supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR), made possible by the Bundesministerium für Wirtschaft und Klimaschutz, under project 50QE2305, an acknowledge funding from the German Science Foundation DFG, via the Collaborative Research Center SFB1491 “Cosmic Interacting Matters - From Source to Signal”. HHi is also supported by a DFG Heisenberg grant (Hi 1495/5-1). HHi, JLvdB, CM, RR, & AD are supported by an ERC Consolidator Grant (No. 770935). MB & PJ are supported by the Polish National Science Center through grant no. 2020/38/E/ST9/00395. MB is also supported by grant no. 2020/39/B/ST9/03494. CH, BS, & ZY acknowledge support from the Max Planck Society and the Alexander von Humboldt Foundation in the framework of the Max Planck-Humboldt Research Award endowed by the Federal Ministry of Education and Research. CH also acknowledges support from the UK Science and Technology Facilities Council (STFC) under grant ST/V000594/1. BJ acknowledges support by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and by STFC Consolidated Grant ST/V000780/1. CM acknowledges support from the Beecroft Trust, and the Spanish Ministry of Science under the grant number PID2021-128338NB-I00. MA is supported by the UK Science and Technology Facilities Council (STFC) under grant number ST/Y002652/1 and the Royal Society under grant numbers RGSR2222268 and ICAR1231094. CG is funded by the MICINN project PID2022-141079NB-C32. BG acknowledges support from the UKRI Stephen Hawking Fellowship (grant reference EP/Y017137/1). HHo & MY acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program with Grant agreement No. 101053992. SJ acknowledges the Ramón y Cajal Fellowship (RYC2022-036431-I) from the Spanish Ministry of Science and the Dennis Sciama Fellowship at the University of Portsmouth. KK acknowledges support from the Royal Society and Imperial College. SSL is receiving funding from the programme “Netzwerke 2021”, an initiative of the Ministry of Culture and Science of the State of North-Rhine Westphalia. LL is supported by the Austrian Science Fund (FWF) [ESP 357-N]. AL acknowledges support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. LM acknowledges the financial contribution from the grant PRIN-MUR 2022 20227RNLY3 “The concordance cosmological model: stress-tests with galaxy clusters” supported by Next Generation EU and from the grant ASI n. 2024-10-HH.0 ‘Attività scientifiche per la missione Euclid – fase E’. LP acknowledges support from the DLR grant 50QE2002. MR acknowledges financial support from the INAF grant 2022. TT acknowledges funding from the Swiss National Science Foundation under the Ambizione project PZ00P2_193352. MvWK acknowledges the support by the UKSA and STFC (grant no. ST/X001075/1). YZ acknowledges the studentship from the UK Science and Technology Facilities Council (STFC). Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme IDs 179.A-2004, 177.A-3016, 177.A-3017, 177.A-3018, 298.A-5015.
@@
Author Contributions: All authors contributed to the development and writing of this paper. The authorship list is given in three groups: the lead authors (AHW, HHi, JLvdB), followed by two alphabetical groups. The first alphabetical group includes those who are key contributors to both the scientific analysis and the data products of this manuscript and release. The second group covers those who have either made a significant contribution to the preparation of data products or to the scientific analyses of KiDS-Legacy.
@@ Appendix A Sample selection function for MICE2
In our clustering redshift analysis with simulated data based on MICE2 we aim to replicate the observational data sets as closely as possible. Since we added spectroscopic data from the DESI EDR and VIPERS PDR-2, we also need to implement their selection functions for MICE2, similar to the procedure already adopted in van den Busch et al. (2020). Where possible, we apply the same selection criteria used for the spectroscopic target selection and use sampling strategies to implement additional selection effects, such as spectroscopic success rates, and to mitigate systematic differences between MICE2 and the observed data sets.
@@ A.1 DESI EDR data
We use a subset (called the LSS catalogues) of the DESI LRG and ELG samples for the KiDS-Legacy clustering redshifts. The target selection for the ELG sample is a simple colour-magnitude cut, which can, in principle, be applied directly to MICE2. The LRG sample, however, and some of the additional selections applied for the construction of the LSS catalogues, depend on observed quantities to which we do not have access to in MICE2. Therefore, we decided to implement the DESI selection function for MICE2 by mostly relying on sampling techniques.
We select the LRG and ELG sample jointly by performing a number of selection steps. First, we split the DESI data and MICE2 into bins of redshift ($\Delta z=0.05$). In each of these bins we compute the expected number of ELGs, LRGs, and (although not utilised) QSOs. Then we take the MICE2 data and randomly draw the appropriate number QSOs by requiring $23.419.5<r<$ 23.4 without any further selections (MICE2 does not contain any QSOs specifically) and discard them. From the remaining MICE2 galaxies we then select the expected number of ELGs by picking the objects with the highest specific star formation rate that fall into the magnitude window $24.120.0<g<$ 24.1. Finally, we draw the expected number of LRGs from those MICE2 galaxies that are not already assigned to either the QSO or ELG sample. We select objects with the highest stellar mass and magnitude $z<21.61$. This procedure ensures that the MICE2 DESI sample has the correct redshift distribution by design. Figure 15 shows the distribution of stellar mass and star-formation rate of galaxies in all of MICE2 and those in our simulated DESI subset. The LRG and ELG subsets are clearly separated in stellar mass, the ELG sample contains mostly objects with low stellar mass but high star-formation rate.
Figure 15: Comparison of star-formation rate (top) and stellar mass (bottom) for the full MICE simulation (gray) and our simulated DESI LRG and ELG samples (blue) as a function of redshift. The data is selected from a $44\leavevmode\nobreak\ \text{deg}^{2}$ patch of MICE2. The lower panel clearly shows the separation of the ELG from the LRG sample, which is selected based on stellar mass.
To verify our new selection function we compare its clustering amplitude $w_{\rm ss}$ with the one we obtain from DESI. We measure the angular correlation between 100 and 1000 kpc and find a good agreement between simulation and data for most redshifts except around the redshifts $0.85$ and $1$ (see Fig. 16). Since we are already selecting objects with low stellar mass in our ELG selection (for which we already expect a lower galaxy bias and therefore a lower clustering amplitude, see e.g. Coil et al. 2017), we speculate that this may be an inherent property of MICE2.
@@ A.2 VIPERS data
We apply the same selection criteria to MICE2.
This colour selection (colour sampling rate; CSR) leads to a completeness that transitions from almost zero to one in the range of $0.60.4<z<$ 0.6. There are two additional effects that need to be factored in to obtain the total completeness of the sample; the target sampling rate (TSR), which is about 50 % on average but has a strong positional dependence due to observational and instrumentational limitations, and the spectroscopic success rate (SSR). The total completeness is a product of these three terms and VIPERS defines a weight to account for this incompleteness as
@@
For our purposes, we choose to not model the positional dependence of the TSR and simply estimate the mean incompleteness weight empirically in the redshift range $1.180.6\leq z<$ 1.18 (see Sect. 2.1), as shown in Fig. 17. When applied to MICE2 together with the VIPERS colour selection, we find that this approach reproduces the redshift distribution $p(z)$ of the VIPERS data set very well. However, we need to apply an additional sparse sampling by 30 % to match the absolute number density found in the data. Similar discrepancies have been reported by van den Busch et al. (2020) when trying to reproduce the selection functions of other high redshift data sets and is most likely explained by systematic differences between MICE2 and the observational data.
Figure 17: Comparison of the VIPERS redshifts distribution (in blue) in the range $1.180.6\leq z<$ 1.18 and MICE2 (in red) after applying the colour/magnitude cuts and the empirical incompleteness sampling, indicated by the total success rate (black line).
Similar to DESI, we verify our new selection function by comparing its clustering amplitude $w_{\rm ss}$ with the one we obtain from VIPERS. We measure on the same scales from 100 and 1000 kpc and find a good agreement between simulation and data, both in the amplitude as well as its uncertainty over the full redshift baseline (Fig. 18).

### v2

@@ Acknowledgements.
AHW and HHi are supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR), made possible by the Bundesministerium für Wirtschaft und Klimaschutz, under project 50QE2305, and acknowledge funding from the German Science Foundation DFG, via the Collaborative Research Center SFB1491 “Cosmic Interacting Matters - From Source to Signal”. HHi is also supported by a DFG Heisenberg grant (Hi 1495/5-1). HHi, JLvdB, CM, RR, & AD are supported by an ERC Consolidator Grant (No. 770935). MB & PJ are supported by the Polish National Science Center through grant no. 2020/38/E/ST9/00395. MB is also supported by grant no. 2020/39/B/ST9/03494. CH, BS, & ZY acknowledge support from the Max Planck Society and the Alexander von Humboldt Foundation in the framework of the Max Planck-Humboldt Research Award endowed by the Federal Ministry of Education and Research. CH also acknowledges support from the UK Science and Technology Facilities Council (STFC) under grant ST/V000594/1. BJ acknowledges support by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and by STFC Consolidated Grant ST/V000780/1. CM acknowledges support from the Beecroft Trust, and the Spanish Ministry of Science under the grant number PID2021-128338NB-I00. MA is supported by the UK Science and Technology Facilities Council (STFC) under grant number ST/Y002652/1 and the Royal Society under grant numbers RGSR2222268 and ICAR1231094. CG is funded by the MICINN project PID2022-141079NB-C32. BG acknowledges support from the UKRI Stephen Hawking Fellowship (grant reference EP/Y017137/1). HHo & MY acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program with Grant agreement No. 101053992. SJ acknowledges the Ramón y Cajal Fellowship (RYC2022-036431-I) from the Spanish Ministry of Science and the Dennis Sciama Fellowship at the University of Portsmouth. KK acknowledges support from the Royal Society and Imperial College. SSL is receiving funding from the programme “Netzwerke 2021”, an initiative of the Ministry of Culture and Science of the State of North-Rhine Westphalia. LL is supported by the Austrian Science Fund (FWF) [ESP 357-N]. AL acknowledges support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. LM acknowledges the financial contribution from the grant PRIN-MUR 2022 20227RNLY3 “The concordance cosmological model: stress-tests with galaxy clusters” supported by Next Generation EU and from the grant ASI n. 2024-10-HH.0 ‘Attività scientifiche per la missione Euclid – fase E’. LP acknowledges support from the DLR grant 50QE2002. MR acknowledges financial support from the INAF grant 2022. TT acknowledges funding from the Swiss National Science Foundation under the Ambizione project PZ00P2_193352. MvWK acknowledges the support by the UKSA and STFC (grant no. ST/X001075/1). YZ acknowledges the studentship from the UK Science and Technology Facilities Council (STFC). Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme IDs 179.A-2004, 177.A-3016, 177.A-3017, 177.A-3018, 298.A-5015.
@@
Author Contributions: All authors contributed to the development and writing of this paper. The authorship list is given in three groups: the lead authors (AHW, HHi, JLvdB), followed by two alphabetical groups. The first alphabetical group includes those who are key contributors to both the scientific analysis and the data products of this manuscript and release. The second group covers those who have made a significant contribution either to the preparation of data products or to the scientific analyses of KiDS-Legacy.
@@ Appendix A Sample selection function for MICE2
In our clustering redshift analysis with simulated data based on MICE2 we aimed to replicate the observational datasets as closely as possible. Since we added spectroscopic data from the DESI EDR and VIPERS PDR-2, we also needed to implement their selection functions for MICE2, similar to the procedure already adopted in van den Busch et al. (2020). Where possible, we applied the same selection criteria used for the spectroscopic target selection and used sampling strategies to implement additional selection effects, such as spectroscopic success rates, and to mitigate systematic differences between MICE2 and the observed datasets.
@@ A.1 DESI EDR data
We used a subset (called the LSS catalogues) of the DESI LRG and ELG samples for the KiDS-Legacy clustering redshifts. The target selection for the ELG sample is a simple colour-magnitude cut, which can, in principle, be applied directly to MICE2. The LRG sample, however, and some of the additional selections applied for the construction of the LSS catalogues, depend on observed quantities to which we do not have access to in MICE2. Therefore, we decided to implement the DESI selection function for MICE2 by mostly relying on sampling techniques.
We selected the LRG and ELG sample jointly by performing a number of selection steps. First, we split the DESI data and MICE2 into bins of redshift ($\Delta z=0.05$). In each of these bins we computed the expected number of ELGs, LRGs, and (although not utilised) quasars (QSOs). Then we randomly draw the appropriate number QSOs from MICE2 by requiring $19.5<r<23.4$ without any further selections (MICE2 does not contain any QSOs specifically) and discarded them. From the remaining MICE2 galaxies we then selected the expected number of ELGs by picking the objects with the highest specific star formation rate that fell in the magnitude window $20.0<g<24.1$. Finally, we drew the expected number of LRGs from those MICE2 galaxies that were not already assigned to either the QSO or ELG sample. We selected objects with the highest stellar mass, and at magnitude $z<21.61$. This procedure ensures that the MICE2 DESI sample has the correct redshift distribution by design. Figure 15 shows the distribution of stellar mass and star-formation rate of galaxies in all of MICE2 and those in our simulated DESI subset. The LRG and ELG subsets are clearly separated in stellar mass, the ELG sample contains mostly objects with low stellar mass but high star-formation rate.
Figure 15: Comparison of the star-formation rate (top) and the stellar mass (bottom) for the full MICE simulation (grey) and our simulated DESI LRG and ELG samples (blue) as a function of redshift. The data is selected from a $44\penalty 10000\ \text{deg}^{2}$ patch of MICE2. The lower panel clearly shows the separation of the ELG from the LRG sample, which is selected based on stellar mass.
To verify our new selection function we compared its clustering amplitude $w_{\rm ss}$ with the one we obtain from DESI. We measured the angular correlation between 100 and 1000 kpc and found good agreement between simulation and data for most redshifts, except around the redshifts $0.85$ and $1$ (see Fig. 16). Since we were already selecting objects with low stellar mass in our ELG selection (for which we already expect a lower galaxy bias and therefore a lower clustering amplitude; see e.g. Coil et al. 2017), we speculate that this may be an inherent property of MICE2.
@@ A.2 VIPERS data
We applied the same selection criteria to MICE2.
This colour selection (colour sampling rate) leads to a completeness that transitions from almost zero to one in the range of $0.4<z<0.6$. There are two additional effects that need to be factored in to obtain the total completeness of the sample; the target sampling rate (TSR), which is about 50 % on average but has a strong positional dependence due to observational and instrumental limitations, and the spectroscopic success rate (SSR). The total completeness is a product of these three terms and VIPERS defines a weight to account for this incompleteness as
@@
For our purposes, we chose to not model the positional dependence of the TSR and simply estimate the mean incompleteness weight empirically in the redshift range $0.6\leq z<1.18$ (see Sect. 2.1), as shown in Fig. 17. When applied to MICE2 together with the VIPERS colour selection, we found that this approach reproduced the redshift distribution $p(z)$ of the VIPERS dataset very well. However, we needed to apply an additional sparse sampling by 30 % to match the absolute number density found in the data. Similar discrepancies have been reported by van den Busch et al. (2020) when trying to reproduce the selection functions of other high redshift datasets and are most likely explained by systematic differences between MICE2 and the observational data.
Figure 17: Comparison of the VIPERS redshifts distribution (in blue) in the range $0.6\leq z<1.18$ and MICE2 (in red) after applying the colour and magnitude cuts, and the empirical incompleteness sampling, indicated by the total success rate (black line).
Similar to DESI, we verified our new selection function by comparing its clustering amplitude $w_{\rm ss}$ with the one we obtained from VIPERS. We measured on the same scales from 100 and 1000 kpc and found a good agreement between simulation and data, both in the amplitude as well as its uncertainty over the full redshift baseline (Fig. 18).